53,670 GDP per Capita (2013) 52,200 46,140 250 26,200 25,920 13,860 0,564 9,940 6,580 5700 2.160 1.570 810 470 In 1950 Americas GDP per capita was appraximately $15,000 (in todays dolilars) How much higher in percentage ts Instructions: Round to the nearest two decimal places (a) Amernicas GDP per capita in 2013 compared to its GDP in 19507 b) How does Chinas and Canadas GDP in 2013 compare to Americas GDP in 19507 )) Chinas in 19507
Answer #1

Answer

A) In 1950 America GDP =$15,000

In 2013 America GDP =$53670

So America % GDP higher = (In 2013 America GDP-In 1950 America GDP)*100/In 1950 America GDP=(53670-15000)*100/15000=257.8%

B)

i)

In 1950 America GDP =$15,000

In 2013 China GDP =$6560

So China % GDP of America GDP in 1950  = (In 2013 China GDP)*100/In 1950 America GDP=6560*100/15000=43.73%

ii)

In 1950 America GDP =$15,000

In 2013 Canada GDP =$52200

So Canada % GDP of America GDP in 1950 = (In 2013 Canada GDP)*100/In 1950 America GDP=52200*100/15000=348%